OpenStudy (anonymous):

4(4k+2)-8(k-2)>3(2k+4)-1

OpenStudy (anonymous):



OpenStudy (unklerhaukus):

expand the brackets

OpenStudy (anonymous):

hwo would i do that? i dont know how to do this ?

OpenStudy (unklerhaukus):

well the first term 4(4k+2) becomes 4*4k+4*2

OpenStudy (unklerhaukus):

can you simplify this ?

OpenStudy (anonymous):

well yeah but i got it wrong on my test

OpenStudy (unklerhaukus):

what do you get when you simplify 4(4k+2)= 4*4k+4*2=

OpenStudy (anonymous):

4(4k+2)=16k+8

OpenStudy (unklerhaukus):

good, now expand the brackets and simplify the text term -8(k-2) = =

OpenStudy (anonymous):

-8(k-2)=16k+8=6k+12-1 right?

OpenStudy (unklerhaukus):

nope -8(k-2) =-8*k+-8*-2 =

OpenStudy (unklerhaukus):

try again

OpenStudy (anonymous):

well sorry but i cnt get it

OpenStudy (anonymous):

for solving the whole problem i got k=2

OpenStudy (anonymous):

Did u get it?

OpenStudy (anonymous):

4(4k+2)-8(k-2)>3(2k+4)-1 Distribute by multiplying 16k+8-8k+16 > 6k+12 -1 Combine like terms 8k+24 > 6k + 11 Subtract 6k from both sides -6k -6k(they cancel each other out) 2k+24> 11 Subtract 24 from both sides -24 -24 2k > -13 Divide both sides by 2 k > -13/2 Answer

OpenStudy (anonymous):

Any q's ? :D

OpenStudy (anonymous):

best response ?

OpenStudy (anonymous):

The symbol with line under it is an equal to so, u would write k >= -13/2

OpenStudy (anonymous):

@cmenendez55 ????? I have to go ...